(10%) Problem 1: A 0475-kg mass suspended from a spring undergoes simple harmonic oscillations with a period of 1.7 s. How much mass, in kilograms, must be added to the object to change the period to 2.05 s? tan() |π|( acos0 sin0 cos0 cotan 1 2 3 0 atan0 acotansinhO coshO tanh0 cotanhO O Degrees O Radians END Submit Hint I give up! Hints: 1% deduction per hint. Hints remaining Feedback:--deduction per feedback.
Answer #1

initial mass of system =m1 =0.475 kg

initial period of oscillation =T1= 1.7 seconds

final period of oscillation required =T2= 2.05 s

let final required mass of system = m2

m2/m1 = sqrt(2.05/1.7)=1.098

m2 = 1.098*0.475=0.5216 kg

additional mass required = 0.5216-0.475=0.0466 kg